A body travels 10m in first three seconds and 15m in the next four seconds. What is the velocity at the end of the 8th second?

Grade:11

1 Answers

Nirmal Singh.
10 years ago

Let's consider initial velocity is u and acceleration is a and 8th sec velocity is v..

we will this equation S=ut+0.5at2 as shown below

10 = u*3 +0.5 *a*3*3 -----1
25 = u*7 + 0.5*a*7*7 ------2
eq 1 => u=(10-4.5*a)/3

Putting above value in eq 2 we get

25 = ((10-4.5*a)*7)/3 + a*24.5
So, a = 5/42 and u =3.15
Now, v = 3.15 + 5/42(8)
So, v=4.1 m/s

Thanks & Regards,

Nirmal Singh

Askiitians Faculty